### v1

# Keyed Chaotic Tensor Transformations for Secure and Attributable Neural Inference

Abstract
This work introduces a novel framework for secure and privacy-preserving neural network inference based on keyed chaotic dynamical transformations. The proposed method applies a deterministic, cryptographically seeded chaotic system to tensors, producing non-invertible, user-specific transformations that enable authenticated inference, tensor-level watermarking, and data attribution. This framework offers a scalable and lightweight alternative to conventional cryptographic techniques, and establishes a new direction for tensor-level security in AI systems.
Note: This is a preliminary preprint released to share the core ideas and invite feedback. Empirical validation and formal security analysis are planned for revisions.
1 Introduction
Modern machine learning systems increasingly operate in settings where user privacy, data security, and attribution are critical. inference pipelines often transmit sensitive user data to hosted models. In these contexts, users must trust that their data are handled securely, and that their contributions are not exposed, misused, or misattributed. Existing approaches to privacy-preserving inference often rely on heavyweight cryptographic methods such as homomorphic encryption [1], and secure multi-party computation [2], or introduce performance trade-offs through techniques like differential privacy [3] and training-time watermarking [4]. These approaches can require significant computational overhead, retraining, or infrastructure changes, limiting their practicality in real-world inference pipelines.
To the best of my knowledge, no existing method offers a lightweight, deterministic approach to tensor-level security that supports both input privacy and contribution traceability, while remaining fully compatible with standard model architectures. Ideally, such a method would act as a cryptographic wrapper for input embeddings, applied either prior to or after inference, and would support a wide range of downstream use cases including authentication, watermarking, attribution, and data contribution auditing.
This work introduces a framework for applying keyed chaotic dynamical transformations to tensors used in machine learning pipelines, enabling privacy-preserving, traceable, and non-invertible manipulation of data representations. Each transformation is derived from a cryptographically seeded dynamical system, represented as a computation graph that operates deterministically on the input. The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are non-invertible, watermarked, and cryptographically bound to a key. These transformations can be applied generally to tensors and are designed to integrate cleanly with existing models, without modifying architecture or requiring retraining.
The main contributions of this work are:
  • A method for applying deterministic chaotic transformations to tensors using keyed graph dynamical systems
  • A demonstration of how these transformations enable authenticated inference, watermarking, and attribution
  • A discussion of downstream applications including secure data contribution, dynamic ML firewalls, and tensor hash functions
This framework bridges ideas from cryptography, dynamical systems, and representation learning, offering a new lens on how models and data can interact securely. Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions. The following sections outline the system design, present empirical results on secure inference and watermark robustness, and explore future directions in secure, chaos-based ML infrastructure.
2 Background and Related Work
2.1 Cryptography
Modern cryptographic systems rely on mathematical transformations that are deterministic, non-invertible, and key-dependent. These properties are used to ensure data confidentiality, integrity, and authenticity across a wide range of communication protocols [5]. While most cryptographic methods operate on discrete data (e.g., byte strings), the key ideas — particularly keyed deterministic transforms and collision-resistant signatures — are relevant when designing mechanisms for secure computation in neural systems.
Two foundational building blocks in this space are symmetric-key encryption and cryptographic hash functions. Symmetric schemes such as AES and HMAC rely on a shared secret key to encrypt or authenticate data, ensuring that only entities with access to the key can reproduce or verify the transformation [6, 7]. These methods are computationally efficient and deterministic, but are designed for bitwise or block-level data and do not preserve semantic or geometric structure — making them incompatible with neural network pipelines where data often lives in high-dimensional continuous tensor spaces.
Cryptographic hash functions such as SHA-256 offer another relevant abstraction: they produce outputs that are deterministic, sensitive to small input changes, and hard to invert or collide [8]. However, they map inputs to fixed-length binary strings and do not support inverse gradients, interpolation, or alignment with downstream learning tasks.
The method introduced in this work draws inspiration from these primitives but is fundamentally distinct: it operates in continuous space, supports tensor-valued inputs, and is task-preserving by design. By using keyed chaotic systems to transform tensors, the method introduces cryptographically inspired properties—such as key-dependent transformation and input sensitivity—while remaining lightweight and compatible with standard model architectures.
2.2 Privacy-Preserving Inference
As machine learning systems become increasingly deployed in real-world, cloud-based, and client-server settings, privacy-preserving inference has emerged as a critical area of research. In many applications, users transmit sensitive data — such as biometrics, medical signals, personal text, or proprietary feature representations — to centralized models for inference. This has raised concerns not only about data exposure, but also about the reusability, traceability, and control of these contributions once they enter the inference pipeline.
A number of approaches have been proposed to mitigate these risks, but each carries substantial trade-offs in performance, infrastructure complexity, or applicability. Techniques such as homomorphic encryption (HE) [1] allow models to perform inference directly on encrypted data, preserving confidentiality without revealing inputs to the model host. However, HE-based approaches are often computationally intensive and have orders of magnitude slower model inference, limiting their deployment to narrow use cases. Similarly, secure multi-party computation (SMPC) distributes computation across parties such that no individual party has full access to the data, but these methods typically require multiple rounds of communication, synchronized computation, and significant coordination overhead [2].
Other approaches focus on statistical privacy. For example, differential privacy (DP) introduces controlled noise into model training or inference pipelines to mask individual contributions [3]. While DP offers formal privacy guarantees, it is stochastic, lossy, and often difficult to apply meaningfully in one-off inference scenarios. Lastly, training-time watermarking techniques embed identifiable signatures into the model’s parameters during training, enabling ownership verification or post hoc misuse detection [4]. However, these methods are typically static, global, and tied to the model itself—making them difficult to extend to per-user or inference-time watermarking, and often requiring access to the model weights.
The approach introduced in this work departs from these paradigms. Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior to or following inference. These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally obfuscated, and task-compatible. While the transformations are reversible by the key holder (e.g., via subtraction or reapplication), they remain opaque to unauthorized parties without knowledge of the underlying system. This enables lightweight, key-authenticated inference, where only correctly transformed inputs yield meaningful results, and offers tensor-level control over privacy, traceability, and secure data contribution—without requiring architectural changes or incurring the computational costs of conventional cryptographic methods.
2.3 Chaotic Dynamical System
Chaotic dynamical systems are deterministic processes that exhibit sensitive dependence on initial conditions, nonlinear evolution, and long-term unpredictability despite being fully governed by known rules. They are widely studied in physics, biology, and nonlinear systems theory, and have seen limited but notable use in machine learning—primarily as components in reservoir computing.
In the context of this work, chaos is used as a security primitive rather than a model architecture. Chaos is associated with the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small changes in initial conditions. The key properties of chaotic systems that make them suitable for this purpose include:
  • Determinism: a fixed chaotic system produces consistent trajectories from identical inputs.
  • Sensitivity to initial conditions: ensures that different inputs (or different users) produce non-overlapping, difficult-to-invert trajectories.
  • Structural complexity: allows for transformation of input tensors in ways that preserve local information but obscure global reversibility.
This work draws inspiration from systems such as coupled map lattices, chaotic attractors, and graph-structured reservoirs, using them to implement transformations that are key-conditioned, lightweight, and difficult to approximate or reverse without full knowledge of the graph topology and dynamical rules. By applying these systems to input or output tensors, it becomes possible to encode user-specific behavior, embed cryptographic structure into data, and control the visibility or traceability of representations within inference pipelines.
3 Keyed Chaotic Transformation Construction
The goal of this method is to construct a graph-based dynamical system whose behavior is key-conditioned and exhibits characteristics of chaos—including sensitivity to initial conditions, deterministic unpredictability, and structural complexity. This system serves as a transformation mechanism for input or output tensors, enabling secure and traceable inference without modifying the model architecture. The construction process consists of sampling a graph topology, selecting a dynamical update rule, evaluating chaotic behavior (e.g., via Lyapunov exponents), and deterministically assembling the system using a cryptographic key. The following subsections outline each stage in detail.
3.1 Graph Topology Sampling
The chaotic system used for transformation is implemented over a graph, where nodes represent internal state variables and edges define the pattern of dynamical interaction. The topology of this graph is sampled deterministically from a cryptographic key, which seeds a pseudorandom number generator. The goal is to produce a fixed but key-specific structure that exhibits rich dynamical behavior when coupled with nonlinear update rules.
Given a cryptographic key $k\in\{0,1\}^{n}$, a pseudorandom seed is derived as
$\text{seed}=\texttt{Hash}(k)\bmod 2^{32},$ (1)
where Hash is a standard cryptographic hash function (e.g., SHA-256), and the result is used to initialize a pseudorandom number generator. This generator is then used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
In this work, topologies are sampled from families such as Erdős–Rényi graphs and Watts–Strogatz small-world graphs, though the framework generalizes to other constructions. For Erdős–Rényi graphs, edges are sampled independently:
$\mathbb{P}\big{(}(i,j)\in E\big{)}=p\quad\text{for all }i\neq j.$ (2)
for a fixed edge probability $p\in(0,1)$ and edge list $E$. Optionally, edge weights $w_{ij}\in[-a,a]$ may be sampled and scaled to control spectral properties. For small-world graphs, a Watts–Strogatz procedure is used, beginning from a $k$-regular ring lattice with each edge rewired independently with probability $\beta$.
Graphs are directed, may be weighted or unweighted, and typically disallow self-loops. The number of nodes is chosen to match the dimensionality of the input tensor segment being transformed (e.g., embedding size or output dimensionality).
Preliminary observations suggest that the structure of the graph strongly influences the resulting dynamics—including stability, memory, and sensitivity to perturbation. For example, graphs with higher connectivity (or spectral radius) may produce more stable or diffusive behavior, while sparse or modular graphs may exhibit more chaotic evolution. However, a full characterization of the relationship between graph topology and dynamical behavior is left as future work and presents a rich area for both empirical and theoretical exploration.
3.2 Dynamical Rule Selection
Each node in the system evolves according to a discrete-time update rule that combines a local nonlinear map with global coupling from neighboring nodes. The resulting dynamical behavior emerges from the interaction between the node-level nonlinearity and the graph structure. After this update, an optional leaky integration step is applied to blend the new state with the previous one.
Let $\mathbf{x}(t)\in\mathbb{R}^{d}$ represent the state of the system at time $t$, and let $A\in\{0,1\}^{d\times d}$ and $W\in\mathbb{R}^{d\times d}$ denote the binary adjacency matrix and edge weight matrix, respectively. The state update proceeds in two stages:
3.2.1 Model-specific update rule:
$\tilde{\mathbf{x}}(t+1)=f\left(\mathbf{x}(t),A,W;\theta\right)$
where $f$ denotes the chosen node-level update rule (e.g., logistic map, Izhikevich, linear), and $\theta$ represents optional rule-specific parameters (e.g., $r$, $\gamma$, thresholds). In most cases, $f$ incorporates both local nonlinear updates and graph-mediated coupling via the weighted adjacency matrix.
3.2.2 Leaky integration (optional):
$\mathbf{x}(t+1)=(1-\alpha)\cdot\mathbf{x}(t)+\alpha\cdot\tilde{\mathbf{x}}(t+1)$
where $\alpha\in[0,1]$ is the leak rate.
In this work, several node-level maps are supported:
  • Logistic map: $f(x_{i})=rx_{i}(1-x_{i})$, where $r\in(3.57,4)$ controls chaotic intensity.
  • Linear map: $f(x_{i})=x_{i}$, for baseline comparison.
  • Izhikevich dynamics: A two-variable spiking neuron model used to explore biologically inspired dynamics.
At runtime, the update rule is selected from a library of registered dynamics, and its parameters are sampled from known regimes that exhibit chaotic behavior. This selection is deterministically conditioned on the cryptographic key, enabling the transformation system to be both modular and reproducible across users and devices.
3.3 Chaos Diagnostics and Characterization
To ensure that the graph dynamical systems constructed using cryptographic keys exhibit chaotic behavior, we perform a series of diagnostics drawn from nonlinear systems theory. Chaos is typically defined by the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small differences in initial conditions. This property is essential for ensuring that similar inputs or neighboring keys result in divergent transformation trajectories, providing both privacy and resistance to inversion.
3.3.1 Lyapunov Exponent Estimation.
To assess whether the system exhibits chaotic behavior, we estimate the largest Lyapunov exponent using a standard two-trajectory method. We initialize two states: $\mathbf{x}_{0}$ and a perturbed version $\hat{\mathbf{x}}_{0}=\mathbf{x}_{0}+\epsilon$, where $\epsilon$ is a small positive scalar. Both trajectories are evolved synchronously under the same dynamical system. At each step, the separation $\delta(t)=\hat{\mathbf{x}}(t)-\mathbf{x}(t)$ is measured, and the Lyapunov exponent is estimated by averaging the log growth rate:
$\lambda\approx\frac{1}{T}\sum_{t=1}^{T}\log\left(\frac{\|\delta(t)\|}{\epsilon}\right)$
To ensure numerical stability, the perturbed state is renormalized after each step:
$\hat{\mathbf{x}}(t)\leftarrow\mathbf{x}(t)+\epsilon\cdot\frac{\delta(t)}{\|\delta(t)\|+\varepsilon}$
where $\varepsilon$ is a small constant to avoid division by zero. A consistently positive value of $\lambda$ indicates exponential divergence and confirms the presence of chaotic dynamics.
3.4 Key-Conditioned System Assembly
The final stage of the construction process assembles the graph dynamical system by combining the sampled topology, selected dynamical rule, and initial state—all deterministically derived from a shared cryptographic key. This key serves as the sole source of entropy for the transformation and ensures that the system is uniquely bound to the key holder.
To maintain predictable chaotic behavior, the method restricts sampling to a curated set of parameter ranges known to yield high sensitivity and non-periodicity. Each candidate configuration includes:
  • A graph family (e.g., Erdős–Rényi, Watts–Strogatz)
  • Associated hyperparameters (e.g., edge probability $p$, rewiring rate $\beta$, node degree $k$)
  • A dynamical rule (e.g., logistic map, Izhikevich system)
  • Rule-specific parameters (e.g., $r$ for logistic map, coupling strength $\gamma$, integration constants)
These configurations are indexed, and a pseudorandom generator seeded from the key selects from this indexed set. This ensures that the resulting system is both deterministic and cryptographically unique, while also avoiding regimes where dynamics are trivial or periodic.
In addition to the structural parameters, the initial state of the system is randomly sampled using the same seed derived from the key. This means both the system’s structure and starting conditions are key-dependent, ensuring that even if two users select the same dynamical configuration, their trajectories will diverge due to differing initial conditions.
The result is a fully assembled, chaotic transformation system that is:
  • Deterministic: reproducible with the key, without requiring key storage
  • Secure: opaque to observers without the key
  • Flexible: compatible with a range of graph and dynamic structures
This construction enables lightweight, cryptographically bound transformations.
3.5 Security Properties
The proposed system provides the following properties:
  • Key-conditioned determinism: Each user’s transformation system is deterministically derived from their cryptographic key, ensuring reproducibility without shared state or explicit key transmission.
  • Input obfuscation: Transformed inputs are non-trivial to interpret without the key, due to both structural scrambling and sensitivity to initialization. Small changes in the key or input yield divergent outputs.
  • Attribution and binding: Only a key holder can produce a given transformation. This allows inference outputs to be linked to specific users or devices, supporting attribution and misuse detection.
  • Opaque trajectories: Without knowledge of the graph topology, weights, and update rule, reverse-engineering the transformation is computationally impractical. The chaotic dynamics ensure exponential divergence from any incorrect reconstruction attempt.
4 Applications
4.1 Key-Authenticated Neural Network Inference
In cloud-hosted or API-based inference systems, it is often desirable to restrict access to only authorized users—both for security and for usage control. Traditional access control mechanisms typically operate at the network or user level—using API keys, authentication tokens, or IP-based restrictions. While these methods can identify the requester, they provide no cryptographic binding between the user and the data submitted to the model. As a result, once access is granted, any input can be submitted, including replayed, forged, or malicious content.
The proposed transformation scheme shifts authentication to the data level by requiring that inputs be correctly transformed using a user-specific, key-conditioned chaotic system. This approach ensures that only users with the correct cryptographic key can produce inputs that yield valid model outputs, effectively binding model access to both the identity and integrity of the data itself. It offers an additional layer of protection against misuse, credential leakage, and unauthorized inference—without requiring changes to the model or encrypted communication channels.
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a sequence of token embeddings. A user constructs a chaotic transformation system from a cryptographic key $k$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$. The input is scrambled client-side via:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
and transmitted to the model. At inference time, the model internally reconstructs the true input by regenerating the same chaotic system and subtracting the mask:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
Only if the correct key is used will this inversion yield a valid input. Any incorrectly scrambled input results in faulty or degraded inference, effectively binding model access to possession of the correct key.
Optionally, the model can also apply a forward scrambling to its outputs using a second chaotic trajectory:
$\tilde{\mathbf{y}}=\mathbf{y}+\mathbf{S}^{\prime}_{k}$
The client then performs the inverse:
$\mathbf{y}=\tilde{\mathbf{y}}-\mathbf{S}^{\prime}_{k}$
This creates a fully authenticated inference pipeline, where both input and output representations are cryptographically tied to the user’s key. This transformation operates independently of network-layer encryption, and can be combined with standard secure transport protocols to protect data in transit.
4.2 Watermarking for Model & User Attribution
In multi-user inference systems, it is often desirable to verify which user or device generated a particular output—especially in contexts involving auditing, content moderation, or abuse prevention. While some systems incorporate output watermarking to identify model-generated content, these are typically model-wide and do not support per-user attribution.
The proposed framework enables per-user output watermarking by applying a deterministic, key-conditioned transformation to the model output. Let $\mathbf{y}\in\mathbb{R}^{d}$ represent the output of a model (e.g., a logit vector, embedding, or generated content representation). A user-specific watermark is applied as:
$\tilde{\mathbf{y}}=\mathbf{y}+\mathbf{S}^{\prime}_{k}$
where $\mathbf{S}^{\prime}_{k}\in\mathbb{R}^{d}$ is a deterministic mask generated from the user’s key using the same chaotic transformation framework. This watermark is embedded server-side before the output is returned to the user or downstream system.
Crucially, this transformation acts as a soft fingerprint: the output remains usable for its original task, but contains a key-specific signal that can later be used to attribute the output to a specific user. If a model output appears in the wild, the server or auditor can regenerate $\mathbf{S}^{\prime}_{k}$ for known keys and match it against the suspicious output to infer its origin.
This watermarking mechanism enables per-user attribution without modifying the model, embedding user IDs, or relying on metadata. It offers a lightweight path to traceable outputs in shared inference environments, content moderation systems, or multi-tenant deployment pipelines.
4.3 Data Privacy
Many inference systems operate on embeddings derived from sensitive user data, such as personal text, biometric signals, or proprietary representations. In these settings, raw embeddings are often transmitted to centralized models for downstream inference—posing a significant privacy risk if intercepted or logged. While full encryption can provide protection, it typically requires secure execution environments or homomorphic computation, which introduces significant overhead.
The proposed transformation scheme offers a lightweight alternative by obfuscating embeddings directly in their native vector space. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of token or patch embeddings. The user applies a chaotic, key-conditioned transformation as follows:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
where $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ is a deterministic mask generated from a graph dynamical system seeded by the cryptographic key $k$. The scrambled embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically obfuscated and difficult to interpret without knowledge of the system.
Importantly, because the transformation is deterministic and reversible with the key, the model hosted on the server can decode the input by regenerating the same mask internally:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
This obfuscation protects the user’s input from inspection in transit or at rest, while still enabling standard inference with minimal computational overhead. Furthermore, reverse engineering the obscured tensors is impractical without access to the key generating the chaotic transformation of the tensor hence securing the users raw data.
4.4 Secure Data Contribution to Shared Models
In collaborative machine learning systems—such as federated learning, crowd-sourced training, or multi-institutional data sharing—users may wish to contribute data to a central model while retaining control over their raw inputs. However, direct data sharing raises concerns about privacy, misattribution, and unauthorized reuse, particularly in domains involving personal or proprietary information.
The proposed transformation framework allows users to submit data obfuscated by a key-conditioned chaotic transformation. Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of input embeddings, and let $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ be the mask generated from the user’s key $k$. The transformed data is computed as:
$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
and contributed to a shared model or data repository. The transformation preserves structure and dimensionality, but obfuscates semantic content, thereby protecting raw data from inspection or unauthorized reuse.
In some scenarios, the server may invert the transformation:
$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
to recover the original data for training. Alternatively, the model could be trained directly on $\tilde{\mathbf{X}}$, treating the transformation as a privacy-preserving encoding layer. However, training directly on obfuscated data may degrade performance depending on the task, and its effectiveness is likely to be application-specific. Further empirical study is needed to determine the conditions under which learning from transformed data remains effective.
5 Discussion
This work introduces a new class of tensor-level transformations based on keyed chaotic dynamical systems, designed to provide input obfuscation, output watermarking, and authenticated inference in modern machine learning pipelines. The system is lightweight, deterministic, and modular—requiring no changes to model architecture.
While the system provides functional security properties such as key-conditioned masking, user attribution, and trajectory-based unpredictability, it does not yet offer formal cryptographic guarantees. In particular, the transformation is reversible by design for authorized users, and may be vulnerable to chosen-input or model inversion attacks in some threat models. As such, it is best viewed as a complementary privacy mechanism—sitting between full encryption and raw data sharing—and may be combined with other techniques (e.g., differential privacy, TEEs, encrypted transport) in high-risk applications.
The choice of dynamical system (e.g., graph topology, update rule, initialization) plays a significant role in determining the transformation’s properties. While this work restricts itself to a curated set of configurations with empirically chaotic behavior, future work could explore learned dynamics, adversarial training of transformation functions, or differentiable chaos-conditioned layers for integration into model backbones.
More broadly, this work suggests that chaotic systems—often studied in the context of memory, control, or biological modeling—can be repurposed as lightweight cryptographic primitives for AI deployment. As the interface between private data and hosted models becomes increasingly important, new methods that offer user-level control, attribution, and secure signaling will be essential. This framework provides one such direction and opens several avenues for further development.
6 Conclusion
This work introduces a novel framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning. By leveraging deterministic chaos seeded from cryptographic keys, the method enables input obfuscation, output watermarking, and authenticated inference without modifying the underlying model. The approach is lightweight, task-compatible, and suitable for a range of real-world applications where privacy, attribution, or secure contribution is essential. As the need for secure AI deployment grows, this method offers a new building block at the interface of dynamical systems, cryptography, and neural computation.
7 Acknowledgements
The author gratefully acknowledges the UKRI PhD studentship at the University of Edinburgh, which supported prior research training and early exploration of related ideas. This work was completed independently during a period of academic leave. The author also acknowledges the use of large language models (LLMs) to support drafting, editing, and idea development during the research and writing process.

### v2

@@
-Keyed Chaotic Tensor Transformations for Secure and Attributable Neural Inference
+Keyed Chaotic Masking: A Functional Privacy Framework for Neural Inference
 Abstract
-This work introduces a novel framework for secure and privacy-preserving neural network inference based on keyed chaotic dynamical transformations.
-The proposed method applies a deterministic, cryptographically seeded chaotic system to tensors, producing non-invertible, user-specific transformations that enable authenticated inference, tensor-level watermarking, and data attribution.
-This framework offers a scalable and lightweight alternative to conventional cryptographic techniques, and establishes a new direction for tensor-level security in AI systems.
-Note: This is a preliminary preprint released to share the core ideas and invite feedback. Empirical validation and formal security analysis are planned for revisions.
+This work introduces a lightweight framework for privacy-preserving neural network inference based on keyed chaotic masking a deterministic, user-specific obfuscation method derived from cryptographically seeded chaotic dynamical systems.
+The approach applies masks to input and output tensors using key-conditioned graph dynamics, enabling authenticated inference, user attribution, and soft output watermarking without modifying model architectures.
+While the underlying chaotic system used to generate each mask is not analytically invertible, the masking operation itself is algebraically reversible by authorized key holders, offering functional privacy without formal cryptographic guarantees.
+Unlike traditional encryption or secure multi-party computation, this method operates in continuous space and imposes minimal computational overhead.
+We describe the construction of the masking system, including graph sampling, dynamical rule selection, and chaos diagnostics.
+Applications include privacy-preserving inference, secure data contribution, and per-user watermarking in shared model pipelines.
+This framework offers a practical and modular building block for user-controlled privacy in modern AI systems.
+Note: This is a preliminary preprint released to share the core ideas and invite early feedback.
+The proposed framework uses cryptographically derived keys (e.g., via ECDH and HKDF) to seed user-specific transformations, and inherits security from these primitives.
+However, the chaotic transformation itself is not a formally proven cryptographic primitive.
+It should be viewed as a functional privacy mechanism, offering deterministic obfuscation and user attribution.
+Future work will explore its robustness under adversarial threat models and its relationship to formal notions of security.
 1 Introduction
@@
 Each transformation is derived from a cryptographically seeded dynamical system, represented as a computation graph that operates deterministically on the input.
-The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are non-invertible, watermarked, and cryptographically bound to a key.
+The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are effectively non-invertible due to chaotic divergence, watermarked, and cryptographically bound to a key.
 These transformations can be applied generally to tensors and are designed to integrate cleanly with existing models, without modifying architecture or requiring retraining.
 The main contributions of this work are:
-  • A method for applying deterministic chaotic transformations to tensors using keyed graph dynamical systems
-  • A demonstration of how these transformations enable authenticated inference, watermarking, and attribution
-  • A discussion of downstream applications including secure data contribution, dynamic ML firewalls, and tensor hash functions
+  • A framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning pipelines, enabling user-specific, deterministic, and lightweight obfuscation of data.
+  • A construction method for cryptographically seeded graph-based dynamical systems, including deterministic graph sampling, node-level chaotic update rules, and Lyapunov-based validation.
+  • A conceptual demonstration of the method’s utility in input obfuscation, output watermarking, and authenticated inference, showing how these transformations support data privacy, user attribution, and secure access.
+  • An exploration of downstream use cases, including privacy-preserving inference, federated data contribution, and soft watermarking for multi-user systems—highlighting the compatibility of the approach with existing model architectures.
 This framework bridges ideas from cryptography, dynamical systems, and representation learning, offering a new lens on how models and data can interact securely.
 Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions.
-The following sections outline the system design, present empirical results on secure inference and watermark robustness, and explore future directions in secure, chaos-based ML infrastructure.
+The following sections outline the system design, propose future experiments on secure inference and watermark robustness, and explore directions for chaos-based ML infrastructure.
 2 Background and Related Work
@@
 Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior to or following inference.
-These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally obfuscated, and task-compatible.
-While the transformations are reversible by the key holder (e.g., via subtraction or reapplication), they remain opaque to unauthorized parties without knowledge of the underlying system.
+These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally transformed, and task-compatible (i.e., preserving shape and structure for downstream model compatibility).
+While the masking operation is algebraically reversible by the key holder (e.g.
+via subtraction), the underlying chaotic system used to generate the mask is not analytically invertible.
+As a result, the overall transformation remains opaque to unauthorized parties.
+The chaotic nature of the transformation ensures exponential sensitivity to small changes in initial state or key, rendering inversion impractical without full system knowledge—even though the transformation is technically deterministic.
 This enables lightweight, key-authenticated inference, where only correctly transformed inputs yield meaningful results, and offers tensor-level control over privacy, traceability, and secure data contribution—without requiring architectural changes or incurring the computational costs of conventional cryptographic methods.
@@
 The goal is to produce a fixed but key-specific structure that exhibits rich dynamical behavior when coupled with nonlinear update rules.
-Given a cryptographic key $k\in\{0,1\}^{n}$, a pseudorandom seed is derived as
-$\text{seed}=\texttt{Hash}(k)\bmod 2^{32},$ (1)
-where Hash is a standard cryptographic hash function (e.g., SHA-256), and the result is used to initialize a pseudorandom number generator.
-This generator is then used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
+Given a cryptographically derived key $k$, a pseudorandom byte stream is generated using a deterministic key derivation function (e.g., HKDF with SHA-256).
+This stream is used to seed a pseudorandom number generator for deterministic graph sampling and system assembly.
+When required for compatibility with legacy APIs, a 32-bit integer seed may be extracted from the first four bytes of the stream.
+The resulting generator is used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
 In this work, topologies are sampled from families such as Erdős–Rényi graphs and Watts–Strogatz small-world graphs, though the framework generalizes to other constructions.
@@
 The result is a fully assembled, chaotic transformation system that is:
-  • Deterministic: reproducible with the key, without requiring key storage
-  • Secure: opaque to observers without the key
-  • Flexible: compatible with a range of graph and dynamic structures
+  • Deterministic: reproducible with the key
+  • Key-obscured: opaque to observers without the key
+  • Composable: the system can be integrated into existing pipelines without modifying model architecture or training objectives
 This construction enables lightweight, cryptographically bound transformations.
-3.5 Security Properties
-The proposed system provides the following properties:
-  • Key-conditioned determinism: Each user’s transformation system is deterministically derived from their cryptographic key, ensuring reproducibility without shared state or explicit key transmission.
-  • Input obfuscation: Transformed inputs are non-trivial to interpret without the key, due to both structural scrambling and sensitivity to initialization.
-Small changes in the key or input yield divergent outputs.
-  • Attribution and binding: Only a key holder can produce a given transformation. This allows inference outputs to be linked to specific users or devices, supporting attribution and misuse detection.
-  • Opaque trajectories: Without knowledge of the graph topology, weights, and update rule, reverse-engineering the transformation is computationally impractical.
-The chaotic dynamics ensure exponential divergence from any incorrect reconstruction attempt.
 4 Applications
+The applications described in this section are conceptual demonstrations based on the properties of the transformation system.
+While they are technically implementable and initial implementation confirms Lyapunov divergence across key configurations; empirical evaluations of inference compatibility, robustness to key error, and resistance to inversion attacks are ongoing.
+In all applications, we assume a shared key is either held by the user or securely negotiated with the server via key exchange (e.g., ECDH).
+The transformation can be applied on either side of the interaction, depending on the privacy model and deployment scenario.
 4.1 Key-Authenticated Neural Network Inference
@@
 A user constructs a chaotic transformation system from a cryptographic key $k$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$.
-The input is scrambled client-side via:
+The input is masked client-side via:
 $\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
@@
 Only if the correct key is used will this inversion yield a valid input.
-Any incorrectly scrambled input results in faulty or degraded inference, effectively binding model access to possession of the correct key.
+Any incorrectly masked input results in faulty or degraded inference, effectively binding model access to possession of the correct key.
 Optionally, the model can also apply a forward scrambling to its outputs using a second chaotic trajectory:
@@
 where $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ is a deterministic mask generated from a graph dynamical system seeded by the cryptographic key $k$.
-The scrambled embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically obfuscated and difficult to interpret without knowledge of the system.
+The masked embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically transformed and difficult to interpret without knowledge of the system.
 Importantly, because the transformation is deterministic and reversible with the key, the model hosted on the server can decode the input by regenerating the same mask internally:
@@
 This obfuscation protects the user’s input from inspection in transit or at rest, while still enabling standard inference with minimal computational overhead.
-Furthermore, reverse engineering the obscured tensors is impractical without access to the key generating the chaotic transformation of the tensor hence securing the users raw data.
+Furthermore, reverse engineering the obscured tensors is impractical without access to the full chaotic transformation, as the dynamics induce exponential divergence with respect to both key and input, making inversion intractable without the key.
 4.4 Secure Data Contribution to Shared Models
@@
 However, direct data sharing raises concerns about privacy, misattribution, and unauthorized reuse, particularly in domains involving personal or proprietary information.
-The proposed transformation framework allows users to submit data obfuscated by a key-conditioned chaotic transformation.
+The proposed transformation framework allows users to submit data masked by a key-conditioned chaotic transformation.
 Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of input embeddings, and let $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ be the mask generated from the user’s key $k$.
@@
 Alternatively, the model could be trained directly on $\tilde{\mathbf{X}}$, treating the transformation as a privacy-preserving encoding layer.
-However, training directly on obfuscated data may degrade performance depending on the task, and its effectiveness is likely to be application-specific.
+However, training directly on transformed data may degrade performance depending on the task, and its effectiveness is likely to be application-specific.
 Further empirical study is needed to determine the conditions under which learning from transformed data remains effective.
@@
 The system is lightweight, deterministic, and modular—requiring no changes to model architecture.
-While the system provides functional security properties such as key-conditioned masking, user attribution, and trajectory-based unpredictability, it does not yet offer formal cryptographic guarantees.
+While the system is designed to offer functional security properties such as key-conditioned masking and trajectory-based unpredictability, these remain to be empirically evaluated in practical deployments.
 In particular, the transformation is reversible by design for authorized users, and may be vulnerable to chosen-input or model inversion attacks in some threat models.
@@
 As the need for secure AI deployment grows, this method offers a new building block at the interface of dynamical systems, cryptography, and neural computation.
+Please note that this preprint version is released to share the concept and invite community feedback.
+Validation experiments and security evaluations are ongoing and will appear in subsequent revisions.
 7 Acknowledgements

### v3

@@
-Keyed Chaotic Masking: A Functional Privacy Framework for Neural Inference
+Keyed Chaotic Dynamics for Privacy-Preserving Neural Inference
 Abstract
-This work introduces a lightweight framework for privacy-preserving neural network inference based on keyed chaotic masking a deterministic, user-specific obfuscation method derived from cryptographically seeded chaotic dynamical systems.
-The approach applies masks to input and output tensors using key-conditioned graph dynamics, enabling authenticated inference, user attribution, and soft output watermarking without modifying model architectures.
-While the underlying chaotic system used to generate each mask is not analytically invertible, the masking operation itself is algebraically reversible by authorized key holders, offering functional privacy without formal cryptographic guarantees.
-Unlike traditional encryption or secure multi-party computation, this method operates in continuous space and imposes minimal computational overhead.
-We describe the construction of the masking system, including graph sampling, dynamical rule selection, and chaos diagnostics.
-Applications include privacy-preserving inference, secure data contribution, and per-user watermarking in shared model pipelines.
-This framework offers a practical and modular building block for user-controlled privacy in modern AI systems.
-Note: This is a preliminary preprint released to share the core ideas and invite early feedback.
-The proposed framework uses cryptographically derived keys (e.g., via ECDH and HKDF) to seed user-specific transformations, and inherits security from these primitives.
-However, the chaotic transformation itself is not a formally proven cryptographic primitive.
-It should be viewed as a functional privacy mechanism, offering deterministic obfuscation and user attribution.
-Future work will explore its robustness under adversarial threat models and its relationship to formal notions of security.
+Neural network inference typically operates on raw input data, increasing the risk of exposure during preprocessing and inference.
+Moreover, neural architectures lack efficient built-in mechanisms for directly authenticating input data.
+This work introduces a novel encryption method for ensuring the security of neural inference.
+By constructing key-conditioned chaotic graph dynamical systems, we enable the encryption and decryption of real-valued tensors within the neural architecture.
+The proposed dynamical systems are particularly suited to encryption due to their sensitivity to initial conditions and their capacity to produce complex, key-dependent nonlinear transformations from compact rules.
+This work establishes a paradigm for securing neural inference and opens new avenues for research on the application of graph dynamical systems in neural network security.
 1 Introduction
-Modern machine learning systems increasingly operate in settings where user privacy, data security, and attribution are critical.
-inference pipelines often transmit sensitive user data to hosted models.
-In these contexts, users must trust that their data are handled securely, and that their contributions are not exposed, misused, or misattributed.
-Existing approaches to privacy-preserving inference often rely on heavyweight cryptographic methods such as homomorphic encryption [1], and secure multi-party computation [2], or introduce performance trade-offs through techniques like differential privacy [3] and training-time watermarking [4].
-These approaches can require significant computational overhead, retraining, or infrastructure changes, limiting their practicality in real-world inference pipelines.
-To the best of my knowledge, no existing method offers a lightweight, deterministic approach to tensor-level security that supports both input privacy and contribution traceability, while remaining fully compatible with standard model architectures.
-Ideally, such a method would act as a cryptographic wrapper for input embeddings, applied either prior to or after inference, and would support a wide range of downstream use cases including authentication, watermarking, attribution, and data contribution auditing.
-This work introduces a framework for applying keyed chaotic dynamical transformations to tensors used in machine learning pipelines, enabling privacy-preserving, traceable, and non-invertible manipulation of data representations.
-Each transformation is derived from a cryptographically seeded dynamical system, represented as a computation graph that operates deterministically on the input.
-The transformation can be configured to either preserve task-relevant structure or intentionally obscure it, depending on the application, resulting in representations that are effectively non-invertible due to chaotic divergence, watermarked, and cryptographically bound to a key.
-These transformations can be applied generally to tensors and are designed to integrate cleanly with existing models, without modifying architecture or requiring retraining.
+Modern machine learning systems increasingly operate in environments where user privacy, data security, and attribution are critical.
+Inference pipelines often transmit sensitive user data to hosted models; in these contexts, users must trust that their data are handled securely, and that their contributions are not exposed, misused, or misattributed.
+Existing approaches to privacy-preserving inference often rely on heavyweight cryptographic methods such as homomorphic encryption [1], and secure multi-party computation [2] which require significant computational overhead or infrastructure changes, limiting their practicality in real-world systems and motivating the need for lightweight, deployable alternatives.
+To address this gap, we seek a practical solution for securing real-valued tensors during inference—one that supports both data privacy and attribution without disrupting existing model architectures.
+Ideally, such a solution would act as a cryptographic wrapper for tensors, applied either prior or post inference, without incurring significant computational overhead.
+This work introduces a framework for constructing chaotic graph dynamical systems and applying them to encrypt and decrypt tensors used in machine learning pipelines, enabling privacy-preserving and traceable manipulation of data representations.
+Each dynamical system is constructed from a cryptographic key and is represented as a computation graph that can readily be incorporated into existing neural architectures.
 The main contributions of this work are:
-  • A framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning pipelines, enabling user-specific, deterministic, and lightweight obfuscation of data.
-  • A construction method for cryptographically seeded graph-based dynamical systems, including deterministic graph sampling, node-level chaotic update rules, and Lyapunov-based validation.
-  • A conceptual demonstration of the method’s utility in input obfuscation, output watermarking, and authenticated inference, showing how these transformations support data privacy, user attribution, and secure access.
-  • An exploration of downstream use cases, including privacy-preserving inference, federated data contribution, and soft watermarking for multi-user systems—highlighting the compatibility of the approach with existing model architectures.
-This framework bridges ideas from cryptography, dynamical systems, and representation learning, offering a new lens on how models and data can interact securely.
+  • A framework for constructing chaotic graph dynamical systems from cryptographic keys.
+  • A demonstration of the application of chaotic graph dynamical systems for encrypting and decrypting tensors with a focus on inference pipelines.
+  • A discussion of the implications of this paradigm for future research directions in privacy preserving machine learning.
+This framework bridges ideas from cryptography and dynamical systems offering a new lens on how models and data can interact securely.
 Unlike traditional cryptographic methods, this approach is architecture-agnostic, low-overhead, and enables per-user control over inference contributions.
-The following sections outline the system design, propose future experiments on secure inference and watermark robustness, and explore directions for chaos-based ML infrastructure.
+The following sections outline the approach, initial validation experiments and directions for future work.
 2 Background and Related Work
 2.1 Cryptography
-Modern cryptographic systems rely on mathematical transformations that are deterministic, non-invertible, and key-dependent.
-These properties are used to ensure data confidentiality, integrity, and authenticity across a wide range of communication protocols [5].
+Modern cryptographic systems rely on key-dependent, invertible transformations that enable secure and authorized access to data.
+These transformations, when combined with additional primitives such as cryptographic hashes and digital signatures, support confidentiality, integrity, and authenticity across a wide range of secure communication protocols [3].
 While most cryptographic methods operate on discrete data (e.g., byte strings), the key ideas — particularly keyed deterministic transforms and collision-resistant signatures — are relevant when designing mechanisms for secure computation in neural systems.
-Two foundational building blocks in this space are symmetric-key encryption and cryptographic hash functions.
+Two foundational building blocks in traditional cryptography are symmetric-key encryption and cryptographic hash functions.
 Symmetric schemes such as AES and HMAC rely on a shared secret key to encrypt or authenticate data, ensuring that only entities with access to the key can reproduce or verify the transformation [4, 5].
-These methods are computationally efficient and deterministic, but are designed for bitwise or block-level data and do not preserve semantic or geometric structure — making them incompatible with neural network pipelines where data often lives in high-dimensional continuous tensor spaces.
+These methods are computationally efficient and deterministic, but are designed for bitwise or block-level data and require quantization or serialization to operate on high-dimensional tensors.
+While invertible, they are not natively compatible with neural network pipelines, which often operate on continuous, real-valued representations during inference.
 Cryptographic hash functions such as SHA-256 offer another relevant abstraction: they produce outputs that are deterministic, sensitive to small input changes, and hard to invert or collide [6].
 However, they map inputs to fixed-length binary strings and do not support inverse gradients, interpolation, or alignment with downstream learning tasks.
-The method introduced in this work draws inspiration from these primitives but is fundamentally distinct: it operates in continuous space, supports tensor-valued inputs, and is task-preserving by design.
-By using keyed chaotic systems to transform tensors, the method introduces cryptographically inspired properties—such as key-dependent transformation and input sensitivity—while remaining lightweight and compatible with standard model architectures.
+The method introduced in this work draws inspiration from these primitives but is fundamentally distinct: it operates in continuous space and supports tensor-valued inputs.
+By using keys to construct graph dynamical systems that encrypt tensors, the method introduces cryptographically inspired properties, such as key-dependent transformation and input sensitivity, while remaining lightweight and compatible with standard model architectures.
 2.2 Privacy-Preserving Inference
 As machine learning systems become increasingly deployed in real-world, cloud-based, and client-server settings, privacy-preserving inference has emerged as a critical area of research.
-In many applications, users transmit sensitive data — such as biometrics, medical signals, personal text, or proprietary feature representations — to centralized models for inference.
+In many applications, users transmit sensitive data, such as biometrics, medical signals, personal text, or proprietary feature representations, to centralized models for inference.
 This has raised concerns not only about data exposure, but also about the reusability, traceability, and control of these contributions once they enter the inference pipeline.
@@
 The approach introduced in this work departs from these paradigms.
-Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior to or following inference.
+Rather than encrypting the model or introducing noise, it applies deterministic, key-conditioned transformations to tensors prior or post inference.
 These transformations are derived from chaotic dynamical systems, producing outputs that are key-specific, structurally transformed, and task-compatible (i.e., preserving shape and structure for downstream model compatibility).
 While the masking operation is algebraically reversible by the key holder (e.g.
-via subtraction), the underlying chaotic system used to generate the mask is not analytically invertible.
+via replaying forward dynamics and performing subtraction), it requires a key as with traditional encryption methods like AES.
 As a result, the overall transformation remains opaque to unauthorized parties.
@@
 2.3 Chaotic Dynamical System
-Chaotic dynamical systems are deterministic processes that exhibit sensitive dependence on initial conditions, nonlinear evolution, and long-term unpredictability despite being fully governed by known rules.
-They are widely studied in physics, biology, and nonlinear systems theory, and have seen limited but notable use in machine learning—primarily as components in reservoir computing.
-In the context of this work, chaos is used as a security primitive rather than a model architecture.
-Chaos is associated with the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small changes in initial conditions.
-The key properties of chaotic systems that make them suitable for this purpose include:
+Chaotic dynamical systems are processes that exhibit sensitive dependence on initial conditions, nonlinear evolution, and long-term unpredictability despite often being fully governed by known rules.
+They are widely studied in physics, biology, and nonlinear systems theory [9], and have seen limited but notable use in machine learning—primarily as components in reservoir computing.
+In the context of this work, chaos is used as a security primitive with analogies drawn between the properties of chaotic dynamics and the properties of traditional encryption algorithms.
+Chaos in this work is identified by the presence of a positive Lyapunov exponent [10], indicating exponential sensitivity to small changes in initial conditions.
+The key properties of chaotic systems that make them suitable as a security primitive include:
   • Determinism: A fixed chaotic system produces consistent trajectories from identical inputs.
-  • Sensitivity to initial conditions: ensures that different inputs (or different users) produce non-overlapping, difficult-to-invert trajectories.
-  • Structural complexity: allows for transformation of input tensors in ways that preserve local information but obscure global reversibility.
-This work draws inspiration from systems such as coupled map lattices, chaotic attractors, and graph-structured reservoirs, using them to implement transformations that are key-conditioned, lightweight, and difficult to approximate or reverse without full knowledge of the graph topology and dynamical rules.
-By applying these systems to input or output tensors, it becomes possible to encode user-specific behavior, embed cryptographic structure into data, and control the visibility or traceability of representations within inference pipelines.
-3 Keyed Chaotic Transformation Construction
-The goal of this method is to construct a graph-based dynamical system whose behavior is key-conditioned and exhibits characteristics of chaos—including sensitivity to initial conditions, deterministic unpredictability, and structural complexity.
-This system serves as a transformation mechanism for input or output tensors, enabling secure and traceable inference without modifying the model architecture.
-The construction process consists of sampling a graph topology, selecting a dynamical update rule, evaluating chaotic behavior (e.g., via Lyapunov exponents), and deterministically assembling the system using a cryptographic key.
-The following subsections outline each stage in detail.
-3.1 Graph Topology Sampling
-The chaotic system used for transformation is implemented over a graph, where nodes represent internal state variables and edges define the pattern of dynamical interaction.
-The topology of this graph is sampled deterministically from a cryptographic key, which seeds a pseudorandom number generator.
-The goal is to produce a fixed but key-specific structure that exhibits rich dynamical behavior when coupled with nonlinear update rules.
-Given a cryptographically derived key $k$, a pseudorandom byte stream is generated using a deterministic key derivation function (e.g., HKDF with SHA-256).
-This stream is used to seed a pseudorandom number generator for deterministic graph sampling and system assembly.
-When required for compatibility with legacy APIs, a 32-bit integer seed may be extracted from the first four bytes of the stream.
-The resulting generator is used to deterministically sample a graph $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
-In this work, topologies are sampled from families such as Erdős–Rényi graphs and Watts–Strogatz small-world graphs, though the framework generalizes to other constructions.
-For Erdős–Rényi graphs, edges are sampled independently:
-$\mathbb{P}\big{(}(i,j)\in E\big{)}=p\quad\text{for all }i\neq j.$ (1)
-for a fixed edge probability $p\in(0,1)$ and edge list $E$.
-Optionally, edge weights $w_{ij}\in[-a,a]$ may be sampled and scaled to control spectral properties.
-For small-world graphs, a Watts–Strogatz procedure is used, beginning from a $k$-regular ring lattice with each edge rewired independently with probability $\beta$.
-Graphs are directed, may be weighted or unweighted, and typically disallow self-loops.
-The number of nodes is chosen to match the dimensionality of the input tensor segment being transformed (e.g., embedding size or output dimensionality).
-Preliminary observations suggest that the structure of the graph strongly influences the resulting dynamics—including stability, memory, and sensitivity to perturbation.
-For example, graphs with higher connectivity (or spectral radius) may produce more stable or diffusive behavior, while sparse or modular graphs may exhibit more chaotic evolution.
-However, a full characterization of the relationship between graph topology and dynamical behavior is left as future work and presents a rich area for both empirical and theoretical exploration.
-3.2 Dynamical Rule Selection
+  • Sensitivity to initial conditions: Small changes in initial state result in vastly different trajectories.
+  • Structural complexity: Chaotic transformations encrypt input tensors into seemingly unstructured forms, making them difficult to interpret or invert without full system knowledge.
+In traditional encryption algorithms such as AES, security is achieved through a series of key-dependent, invertible transformations—typically involving substitution-permutation networks and multiple rounds of mixing operations.
+While structurally different, the chaotic transformations used in this work serve a comparable role: they are deterministic, key-conditioned, and designed to obscure the relationship between input and output tensors.
+Like AES, small changes to the key or input yield drastically different outputs, ensuring unpredictability and resistance to unauthorized inversion.
+However, unlike AES, our transformations operate in continuous space, are differentiable, and integrate seamlessly into existing machine learning pipelines—enabling lightweight, task-compatible obfuscation without requiring model modification or cryptographic decoding layers.
+3 Keyed Chaotic Dynamics
+Figure 1: Overview of the keyed chaotic masking and unmasking process.
+A cryptographic key is used to instantiate a chaotic graph dynamical system, which produces a deterministic mask matrix $S$.
+During encryption (top path), this mask is element-wise added to the input tensor X, producing a masked tensor $\tilde{X}=X+S$.
+During decryption (bottom path), the same mask is regenerated using the shared key and subtracted from $\tilde{X}$ to recover the original input: $X=\tilde{X}-S$.
+The central graph diagram represents the chaotic system generated from the key, and the matrices represent tensor-valued data.
+3.1 Key Hashing and Seed Derivation
+The chaotic graph dynamical system used in this work is deterministically instantiated from a cryptographic key through a structured key schedule and parameter sampling process.
+We assume the presence of a secure key exchange protocol, such as Elliptic Curve Diffie-Hellman (ECDH), to derive a shared secret key $k$.
+To prevent reuse of masks across different inputs or sessions, we also introduce a public nonce $n$, which is combined with the key during subkey derivation.
+To securely and modularly configure the chaotic system from this key, we adopt a domain-specific key schedule inspired by symmetric encryption algorithms such as AES.
+We start by deriving multiple subkeys from the master key k by hashing together fixed domain labels and the nonce with the key:
+[…]
+[…]
+[…]
+Each resulting subkey is then used to seed its corresponding sampling procedure: graph topology construction, chaotic map parameter selection, and initial condition generation.
+This separation improves modularity, ensures reproducibility, and avoids parameter entanglement between system components.
+3.2 Constructing Chaotic Graph Dynamical Systems
+The goal of this section is to construct a graph dynamical system whose behavior is key-conditioned and exhibits characteristics of chaos.
+This system serves as the foundation for a transformation mechanism that operates on input or output tensors, enabling secure and traceable inference without modifying neural network architectures.
+3.2.1 Graph Sampling
+The chaotic system used for transformation is implemented over a graph, where nodes represent internal state variables and edges determine the coupling between node dynamics.
+The topology of this graph is sampled deterministically from a cryptographic key, producing a fixed but key-specific structure that supports rich dynamical behavior when combined with nonlinear update rules.
+Given a cryptographically derived subkey $k_{\text{graph}}$, obtained via SHA-256 of a labeled key expansion step, we treat it as a high-entropy seed for initializing a pseudorandom number generator (PRNG).
+This ensures deterministic graph sampling while maintaining strong key-dependent variation.
+If needed for compatibility with legacy APIs, a 32-bit integer seed can be extracted from the first four bytes of $k_{\text{graph}}$.
+The resulting PRNG instance is then used to construct the system’s graph topology $G=(V,E)$, where $|V|=d$ matches the dimensionality of the tensor being transformed.
+In practice we represent the graph with a single matrix $W\in\mathbb{R}^{d\times d}$ representing the coupling strengths between nodes.
+In this work, adjacency matrices $A$ representing graph topologies are first deterministically sampled from classical random graph families, including Erdős–Rényi (ER) and Watts–Strogatz (WS) small-world models.
+For ER graphs, edges are sampled independently with fixed probability $p$, while small-world graphs are generated using a $k$-regular lattice with random rewiring governed by parameter $\beta$.
+Having sampled an adjacency matrix $A$, edge weights which define coupling strengths between nodes are sampled.
+In the simplest case, a weight matrix $W$ is drawn from a uniform distribution over a fixed interval and subsequently the Hadamard product is taken between this matrix and the adjacency matrix as follows:
+$W\leftarrow W\odot A$ (2)
+As with the sampling of adjacency matrices, the current framework generalizes to arbitrary weight sampling distributions, enabling finer control over spectral properties and interaction strength.
+The resulting key-specific graph serves as the backbone of the chaotic system, defining how local dynamics interact and evolve over time.
+3.2.2 Node Dynamics
 Each node in the system evolves according to a discrete-time update rule that combines a local nonlinear map with global coupling from neighboring nodes.
 The resulting dynamical behavior emerges from the interaction between the node-level nonlinearity and the graph structure.
-After this update, an optional leaky integration step is applied to blend the new state with the previous one.
-Let $\mathbf{x}(t)\in\mathbb{R}^{d}$ represent the state of the system at time $t$, and let $A\in\{0,1\}^{d\times d}$ and $W\in\mathbb{R}^{d\times d}$ denote the binary adjacency matrix and edge weight matrix, respectively.
-The state update proceeds in two stages:
-3.2.1 Model-specific update rule:
-$\tilde{\mathbf{x}}(t+1)=f\left(\mathbf{x}(t),A,W;\theta\right)$
-where $f$ denotes the chosen node-level update rule (e.g., logistic map, Izhikevich, linear), and $\theta$ represents optional rule-specific parameters (e.g., $r$, $\gamma$, thresholds).
-In most cases, $f$ incorporates both local nonlinear updates and graph-mediated coupling via the weighted adjacency matrix.
-3.2.2 Leaky integration (optional):
-$\mathbf{x}(t+1)=(1-\alpha)\cdot\mathbf{x}(t)+\alpha\cdot\tilde{\mathbf{x}}(t+1)$
-where $\alpha\in[0,1]$ is the leak rate.
-In this work, several node-level maps are supported:
-  • Logistic map: $f(x_{i})=rx_{i}(1-x_{i})$, where $r\in(3.57,4)$ controls chaotic intensity.
-  • Linear map: $f(x_{i})=x_{i}$, for baseline comparison.
-  • Izhikevich dynamics: A two-variable spiking neuron model used to explore biologically inspired dynamics.
-At runtime, the update rule is selected from a library of registered dynamics, and its parameters are sampled from known regimes that exhibit chaotic behavior.
+Let $\mathbf{x}(t)\in\mathbb{R}^{d}$ represent the state of the system at time $t$, and let $W\in\mathbb{R}^{d\times d}$ denote the sampled weight matrix representing the graph.
+The state update proceeds according to:
+$\tilde{\mathbf{x}}(t+1)=f(\mathbf{x}(t),W;\phi)=f_{\text{local}}(\mathbf{x}(t);\phi)+f_{\text{coupled}}(\mathbf{x}(t),W;\phi)$ (3)
+where $f$ denotes the chosen node-level update dynamics, and $\phi$ represents dynamics-specific parameters.
+In most cases, $f$ incorporates both local nonlinear updates and graph-mediated coupling via the weighted matrix.
+In this work, several candidates for local nonlinear updates are used, including:
+Map Name Update Rule Output State Parameter
+Logistic map $x_{i+1}=rx_{i}(1-x_{i})$ $x_{i}\in(0,1)$
+Tent map $x_{i+1}=\begin{cases}\frac{x_{i}}{\mu}&x_{i}<\mu\\ \frac{1-x_{i}}{1-\mu}&x_{i}\geq\mu\end{cases}$ $x_{i}\in(0,1)$
+Baker’s map $(x_{i+1},y_{i+1})=\begin{cases}(sx_{i},y_{i})&x_{i}<s\\ ((1-s)x_{i}+s,(1-s)y_{i}+s)&x_{i}\geq s\end{cases}$ $x_{i}\in[0,1]$
+[…]
+Arnold’s cat map $\begin{bmatrix}x_{i+1}\\ y_{i+1}\end{bmatrix}=\begin{bmatrix}1&1\\ 1&2\end{bmatrix}\begin{bmatrix}x_{i}\\ y_{i}\end{bmatrix}\mod 1$ $x_{i}\in[0,1]$
+Table 1: Summary of candidate node-level chaotic maps used in this work.
+Different methods of coupling between nodes can be applied to each of the local nonlinear update rules. In this work we choose diffusive coupling between nodes which is defined as:
+$\tilde{x}_{i}(t+1)=f_{\text{local}}(x_{i}(t);\phi)+\sum_{j=1}^{d}W_{ij}\left(f_{\text{local}}(x_{j}(t);\phi)-f_{\text{local}}(x_{i}(t);\phi)\right)$ (4)
+At runtime, the update rule is selected from a library of registered dynamics, and its parameters are sampled from ranges known to result in bounded chaotic dynamics.
 This selection is deterministically conditioned on the cryptographic key, enabling the transformation system to be both modular and reproducible across users and devices.
-3.3 Chaos Diagnostics and Characterization
+3.2.3 Chaotic Dynamics Diagnostics
 To ensure that the graph dynamical systems constructed using cryptographic keys exhibit chaotic behavior, we perform a series of diagnostics drawn from nonlinear systems theory.
 Chaos is typically defined by the presence of a positive Lyapunov exponent, indicating exponential sensitivity to small differences in initial conditions.
-This property is essential for ensuring that similar inputs or neighboring keys result in divergent transformation trajectories, providing both privacy and resistance to inversion.
-3.3.1 Lyapunov Exponent Estimation.
+This property is essential for ensuring that similar inputs or neighboring keys result in divergent transformation trajectories an important property for ensuring the security of this encrpytion method.
 To assess whether the system exhibits chaotic behavior, we estimate the largest Lyapunov exponent using a standard two-trajectory method.
@@
 where $\varepsilon$ is a small constant to avoid division by zero. A consistently positive value of $\lambda$ indicates exponential divergence and confirms the presence of chaotic dynamics.
-3.4 Key-Conditioned System Assembly
-The final stage of the construction process assembles the graph dynamical system by combining the sampled topology, selected dynamical rule, and initial state—all deterministically derived from a shared cryptographic key.
-This key serves as the sole source of entropy for the transformation and ensures that the system is uniquely bound to the key holder.
-To maintain predictable chaotic behavior, the method restricts sampling to a curated set of parameter ranges known to yield high sensitivity and non-periodicity. Each candidate configuration includes:
-  • A graph family (e.g., Erdős–Rényi, Watts–Strogatz)
-  • Associated hyperparameters (e.g., edge probability $p$, rewiring rate $\beta$, node degree $k$)
-  • A dynamical rule (e.g., logistic map, Izhikevich system)
-  • Rule-specific parameters (e.g., $r$ for logistic map, coupling strength $\gamma$, integration constants)
-These configurations are indexed, and a pseudorandom generator seeded from the key selects from this indexed set.
-This ensures that the resulting system is both deterministic and cryptographically unique, while also avoiding regimes where dynamics are trivial or periodic.
-In addition to the structural parameters, the initial state of the system is randomly sampled using the same seed derived from the key.
-This means both the system’s structure and starting conditions are key-dependent, ensuring that even if two users select the same dynamical configuration, their trajectories will diverge due to differing initial conditions.
-The result is a fully assembled, chaotic transformation system that is:
-  • Deterministic: reproducible with the key
-  • Key-obscured: opaque to observers without the key
-  • Composable: the system can be integrated into existing pipelines without modifying model architecture or training objectives
-This construction enables lightweight, cryptographically bound transformations.
+3.3 Encryption Algorithm
+This section consolidates the components described above into a complete encryption and decryption procedure.
+Given a cryptographic key, a deterministic chaotic graph dynamical system is instantiated by deriving subkeys and sampling graph structure and node dynamics.
+Forward dynamics is then simulated to produce a structured mask that is applied to the input tensor via element-wise addition in the case of encryption and subtraction in the case of decryption.
+The resulting process is lightweight, reproducible, and key-dependent—offering functional privacy, it can also be incorporated directly into existing neural architectures as a layer without requiring retraining of the model architecture.
+The full algorithm is outlined below:
+Input: Input tensor $X$, secret key $k$, public nonce $n$
+Output: Masked tensor $\tilde{X}$ or recovered tensor $X$
+Derive subkeys $k_{\text{graph}},k_{\text{params}},k_{\text{init}},k_{\text{noise}}$ using SHA-256:;
+foreach label $i\in\{\texttt{graph},\texttt{params},\texttt{init},\texttt{noise}\}$ do
+       $k_{i}\leftarrow\text{SHA-256}(\texttt{label}_{i}\,\|\,k\,\|\,n)$;
+Construct chaotic graph dynamical system:;
+Sample graph topology $G$ from $k_{\text{graph}}$;
+Configure parameters and initialize state from $k_{\text{params}},k_{\text{init}}$;
+Seed a pseudorandom number generator with $k_{\text{noise}}$ to generate deterministic noise;
+for each timestep $t=1$ to $T$ do
+       Apply node updates using local dynamics and diffusive coupling;
+       Inject deterministic noise into each node state at timestep $t$;
+Aggregate final node states into mask matrix $S\in\mathbb{R}^{n\times d}$;
+Encryption: $\tilde{X}\leftarrow X+S$;
+Decryption: $X\leftarrow\tilde{X}-S$;
+Algorithm 1 Keyed Chaotic Dynamics (Encryption and Decryption with Deterministic Noise)
 4 Applications
-The applications described in this section are conceptual demonstrations based on the properties of the transformation system.
-While they are technically implementable and initial implementation confirms Lyapunov divergence across key configurations; empirical evaluations of inference compatibility, robustness to key error, and resistance to inversion attacks are ongoing.
-In all applications, we assume a shared key is either held by the user or securely negotiated with the server via key exchange (e.g., ECDH).
-The transformation can be applied on either side of the interaction, depending on the privacy model and deployment scenario.
+This section outlines practical use cases enabled by the proposed transformation framework.
+While experimental validation is ongoing, the design properties of the system—key-conditioned masking, deterministic reversibility, and compatibility with real-valued tensor representations—make it suitable for deployment in a range of privacy-sensitive machine learning settings.
+In all cases, we assume a shared secret key $k$ is securely negotiated between parties (e.g., via ECDH) and that a public nonce $n$ is used to ensure per-input uniqueness.
 4.1 Key-Authenticated Neural Network Inference
@@
 Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be a sequence of token embeddings.
-A user constructs a chaotic transformation system from a cryptographic key $k$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$.
-The input is masked client-side via:
+A user constructs a chaotic graph dynamical system from a cryptographic key $k$ and nonce $n$, and generates a deterministic mask $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$.
+The input is encrypted client-side via:
 $\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
@@
 Any incorrectly masked input results in faulty or degraded inference, effectively binding model access to possession of the correct key.
-Optionally, the model can also apply a forward scrambling to its outputs using a second chaotic trajectory:
+Optionally, the model can also encrypt its outputs using a second chaotic trajectory:
 $\tilde{\mathbf{y}}=\mathbf{y}+\mathbf{S}^{\prime}_{k}$
@@
 This creates a fully authenticated inference pipeline, where both input and output representations are cryptographically tied to the user’s key.
-This transformation operates independently of network-layer encryption, and can be combined with standard secure transport protocols to protect data in transit.
-4.2 Watermarking for Model & User Attribution
-In multi-user inference systems, it is often desirable to verify which user or device generated a particular output—especially in contexts involving auditing, content moderation, or abuse prevention.
-While some systems incorporate output watermarking to identify model-generated content, these are typically model-wide and do not support per-user attribution.
-The proposed framework enables per-user output watermarking by applying a deterministic, key-conditioned transformation to the model output.
-Let $\mathbf{y}\in\mathbb{R}^{d}$ represent the output of a model (e.g., a logit vector, embedding, or generated content representation).
-A user-specific watermark is applied as:
-$\tilde{\mathbf{y}}=\mathbf{y}+\mathbf{S}^{\prime}_{k}$
-where $\mathbf{S}^{\prime}_{k}\in\mathbb{R}^{d}$ is a deterministic mask generated from the user’s key using the same chaotic transformation framework.
-This watermark is embedded server-side before the output is returned to the user or downstream system.
-Crucially, this transformation acts as a soft fingerprint: the output remains usable for its original task, but contains a key-specific signal that can later be used to attribute the output to a specific user.
-If a model output appears in the wild, the server or auditor can regenerate $\mathbf{S}^{\prime}_{k}$ for known keys and match it against the suspicious output to infer its origin.
-This watermarking mechanism enables per-user attribution without modifying the model, embedding user IDs, or relying on metadata.
-It offers a lightweight path to traceable outputs in shared inference environments, content moderation systems, or multi-tenant deployment pipelines.
-4.3 Data Privacy
-Many inference systems operate on embeddings derived from sensitive user data, such as personal text, biometric signals, or proprietary representations.
-In these settings, raw embeddings are often transmitted to centralized models for downstream inference—posing a significant privacy risk if intercepted or logged.
-While full encryption can provide protection, it typically requires secure execution environments or homomorphic computation, which introduces significant overhead.
-The proposed transformation scheme offers a lightweight alternative by obfuscating embeddings directly in their native vector space.
-Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of token or patch embeddings.
-The user applies a chaotic, key-conditioned transformation as follows:
-$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
-where $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ is a deterministic mask generated from a graph dynamical system seeded by the cryptographic key $k$.
-The masked embeddings $\tilde{\mathbf{X}}$ retain the structure and dimensions required for downstream tasks, but are semantically transformed and difficult to interpret without knowledge of the system.
-Importantly, because the transformation is deterministic and reversible with the key, the model hosted on the server can decode the input by regenerating the same mask internally:
-$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
-This obfuscation protects the user’s input from inspection in transit or at rest, while still enabling standard inference with minimal computational overhead.
-Furthermore, reverse engineering the obscured tensors is impractical without access to the full chaotic transformation, as the dynamics induce exponential divergence with respect to both key and input, making inversion intractable without the key.
-4.4 Secure Data Contribution to Shared Models
-In collaborative machine learning systems—such as federated learning, crowd-sourced training, or multi-institutional data sharing—users may wish to contribute data to a central model while retaining control over their raw inputs.
-However, direct data sharing raises concerns about privacy, misattribution, and unauthorized reuse, particularly in domains involving personal or proprietary information.
-The proposed transformation framework allows users to submit data masked by a key-conditioned chaotic transformation.
-Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ represent a batch of input embeddings, and let $\mathbf{S}_{k}\in\mathbb{R}^{n\times d}$ be the mask generated from the user’s key $k$.
-The transformed data is computed as:
-$\tilde{\mathbf{X}}=\mathbf{X}+\mathbf{S}_{k}$
-and contributed to a shared model or data repository.
-The transformation preserves structure and dimensionality, but obfuscates semantic content, thereby protecting raw data from inspection or unauthorized reuse.
-In some scenarios, the server may invert the transformation:
-$\mathbf{X}=\tilde{\mathbf{X}}-\mathbf{S}_{k}$
-to recover the original data for training.
-Alternatively, the model could be trained directly on $\tilde{\mathbf{X}}$, treating the transformation as a privacy-preserving encoding layer.
-However, training directly on transformed data may degrade performance depending on the task, and its effectiveness is likely to be application-specific.
-Further empirical study is needed to determine the conditions under which learning from transformed data remains effective.
+This transformation operates independently of network-layer encryption, and can be combined with standard secure transport protocols to protect both data in transit and storage.
+4.2 Privacy-Preserving Data Sharing and Inference
+In both inference and collaborative learning settings, real-valued embeddings may contain sensitive information (e.g., personal text, biometrics, or proprietary features).
+The proposed transformation enables users to encrypt these representations locally using a key-conditioned chaotic system before transmitting them to a server or aggregator.
+The encrypted input $\tilde{X}=X+S_{k}$ retains the structure and shape required for downstream tasks.
+Because the transformation is deterministic and reversible with the key and nonce, the server can reconstruct the original input for inference or training, this is especially relevant in TEE environments where the computational graph representing the encryption/decryption algorithm is not easily interpreted.
+This enables secure, per-user control over data contributions—offering lightweight protection against inspection, misuse, or unauthorized reuse.
+5 Security Considerations
+5.1 Threat Model
+This framework is designed as a lightweight privacy mechanism for inference-time data obfuscation and input authentication. It aims to:
+  • Prevent unauthorized users from generating valid inputs to a model without access to a cryptographic key.
+  • Obfuscate real-valued inputs during transit or storage without modifying model architecture.
+  • Enable reversible masking for authorized parties in possession of the key.
+We assume a standard threat model where:
+  • The shared key $k$ is negotiated securely (e.g., via Elliptic Curve Diffie-Hellman).
+  • The attacker has access to the masked tensor $\tilde{X}$, but not the key or the internal configuration of the mask-generating chaotic system.
+  • The system is not used in repeated or low-entropy input settings without additional safeguards.
+5.2 Security Properties
+The proposed transformation offers the following functional guarantees:
+  • Key- and nonce-conditioned obfuscation: The transformation from $X$ to $\tilde{X}$ is deterministic and reversible only with the correct cryptographic key $k$ and associated nonce $n$.
+This ensures unique masks per input and prevents mask reuse attacks, even when the same key is reused across queries.
+  • Input authentication: Only users in possession of the correct key and nonce can produce valid masked inputs that result in successful model inference, effectively binding access to possession of the correct credentials.
+  • Nonlinear unpredictability: The chaotic update rule introduces exponential sensitivity to both key and input perturbations.
+The injection of deterministic, key-derived noise at each timestep further increases trajectory divergence, degrading the feasibility of surrogate inversion or pattern learning.
+  • Model compatibility: The transformation operates entirely in continuous space and is represented by a computation graph.
+It integrates seamlessly into existing deep learning pipelines and can be implemented efficiently on hardware accelerators such as GPUs using CUDA kernels or fused operations, enabling low-latency deployment even in real-time inference settings.
+5.2.1 Implementation Considerations for Secure Deployment
+The transformation can be efficiently implemented as a fused CUDA kernel to minimize memory access overhead and latency.
+However, this design choice introduces additional considerations: GPU memory and execution contexts are shared across processes, and on systems lacking strict memory isolation, adversaries with privileged access may attempt to inspect device memory or leverage side-channel leakage (e.g., timing or power-based signals) to infer key-conditioned behavior.
+While such attacks require elevated privileges and are not typically feasible in hosted model settings, developers deploying this method in untrusted or multi-tenant environments should consider GPU sandboxing, memory zeroing, and kernel isolation practices.
+Integrating the transformation within a trusted execution environment (TEE) or extending it with secure enclaving techniques may offer additional protection against kernel-level inspection.
+5.3 Limitations and Attack Vectors
+This transformation is best viewed as a functional privacy primitive for continuous tensor data—analogous in structure to symmetric encryption schemes such as AES, but adapted for real-valued, neural architectures.
+While it does not aim to replace formal cryptographic methods in settings requiring ciphertext indistinguishability, it offers a complementary security protocol for scenarios involving neural inference.
+Potential vulnerabilities include:
+  • Model inversion attacks (unlikely): An adversary with full access to masked inputs and model outputs could theoretically attempt to learn a surrogate inverse mapping.
+However, the transformation is highly nonlinear, key-conditioned, and exhibits exponential sensitivity to perturbations via chaotic dynamics.
+Additionally, the injection of deterministic, per-timestep noise renders the resulting mask trajectories resistant to approximation, making such attacks computationally infeasible in practice.
+  • Key compromise: The transformation is fully reversible by design.
+If the cryptographic key and associated nonce are leaked or mishandled, the original input can be trivially recovered by an adversary.
+Standard key management practices (e.g., ephemeral session keys, ECDH negotiation) should be used to mitigate this risk.
+  • Lack of formal obfuscation guarantees: The transformation depending choice of dynamics applies a bijective, key-conditioned nonlinear mapping over continuous, real-valued tensors.
+While the resulting representations are semantically obfuscated and reversible only by authorized parties, the system does not rely on finite-field cryptographic primitives and does not satisfy formal definitions of indistinguishability.
+As such, it should be viewed as a structured, functional privacy mechanism rather than a substitute for conventional encryption in adversarial threat models.
+5.4 Integration with Secure Execution
+This system is intended as a complementary privacy layer. In high-risk deployments, it can be combined with:
+  • Standard transport-layer encryption (e.g., TLS)
+  • Trusted Execution Environments
+  • Ephemeral key derivation or key rotation schemes
+  • Obfuscation techniques such as dropout, quantization, or learned masking
+5.5 Future Directions
+Future work includes:
+  • Formal analysis of the system’s resistance to statistical and adversarial attacks.
+  • Investigation of dynamic or learned chaotic systems for adaptive masking.
+  • Exploration of surrogate model defenses and integration into split-learning architectures.
+This framework is best viewed as a flexible privacy primitive that enables data-level masking and authentication with minimal architectural assumptions.
+It is particularly useful in scenarios where full encryption is infeasible or unnecessary, but lightweight, key-bound transformations are desirable.
 6 Discussion
-This work introduces a new class of tensor-level transformations based on keyed chaotic dynamical systems, designed to provide input obfuscation, output watermarking, and authenticated inference in modern machine learning pipelines.
-The system is lightweight, deterministic, and modular—requiring no changes to model architecture.
-While the system is designed to offer functional security properties such as key-conditioned masking and trajectory-based unpredictability, these remain to be empirically evaluated in practical deployments.
-In particular, the transformation is reversible by design for authorized users, and may be vulnerable to chosen-input or model inversion attacks in some threat models.
-As such, it is best viewed as a complementary privacy mechanism—sitting between full encryption and raw data sharing—and may be combined with other techniques (e.g., differential privacy, TEEs, encrypted transport) in high-risk applications.
-The choice of dynamical system (e.g., graph topology, update rule, initialization) plays a significant role in determining the transformation’s properties.
-While this work restricts itself to a curated set of configurations with empirically chaotic behavior, future work could explore learned dynamics, adversarial training of transformation functions, or differentiable chaos-conditioned layers for integration into model backbones.
-More broadly, this work suggests that chaotic systems—often studied in the context of memory, control, or biological modeling—can be repurposed as lightweight cryptographic primitives for AI deployment.
-As the interface between private data and hosted models becomes increasingly important, new methods that offer user-level control, attribution, and secure signaling will be essential.
-This framework provides one such direction and opens several avenues for further development.
+This work introduces a new class of tensor-level transformations based on keyed chaotic dynamical systems, designed to apply key-conditioned, reversible obfuscation to real-valued tensors within modern machine learning pipelines.
+The system is lightweight, deterministic, and modular—requiring no changes to model architecture and operating natively in continuous space.
+It is amenable to efficient GPU execution, including fused CUDA kernel implementations for real-time inference masking.
+The transformation provides functional security guarantees such as cryptographic key- and nonce-conditioned masking, semantic obfuscation, and trajectory-level unpredictability.
+It supports both input authentication and reversible transformation, enabling inference-time privacy and secure data contribution without disrupting model compatibility.
+While the system is reversible by design for authorized users, its behavior in adversarial settings remains to be evaluated.
+In particular, the method does not rely on formal encryption primitives over finite fields and may be vulnerable to misuse if keys are leaked or mask reuse occurs without appropriate nonce management.
+As such, the transformation is best viewed as a complementary privacy layer—positioned between full encryption and raw data sharing—and can be combined with techniques such as differential privacy, trusted execution environments, or encrypted transport protocols in high-assurance contexts.
+The design of the chaotic system (e.g., graph topology, local map, coupling strategy) plays a critical role in shaping both the security and representational properties of the transformation.
+This work restricts itself to a curated set of update rules with empirically verified chaotic behavior, but future directions include optimizing or learning system configurations from data, developing differentiable chaos layers, or adversarially training transformation functions to withstand surrogate model attacks.
+More broadly, this work suggests that chaotic dynamical systems—traditionally studied in the context of memory, control, or biological computation—can be reimagined as lightweight, cryptographically inspired building blocks for AI systems.
+As machine learning continues to interface with sensitive and user-controlled data, techniques that support secure signaling, data attribution, and behavioral integrity will be increasingly essential.
+This framework provides one such path forward, and opens several directions for continued exploration.
 7 Conclusion
 This work introduces a novel framework for applying keyed chaotic dynamical systems to tensor transformations in machine learning.
-By leveraging deterministic chaos seeded from cryptographic keys, the method enables input obfuscation, output watermarking, and authenticated inference without modifying the underlying model.
-The approach is lightweight, task-compatible, and suitable for a range of real-world applications where privacy, attribution, or secure contribution is essential.
-As the need for secure AI deployment grows, this method offers a new building block at the interface of dynamical systems, cryptography, and neural computation.
-Please note that this preprint version is released to share the concept and invite community feedback.
-Validation experiments and security evaluations are ongoing and will appear in subsequent revisions.
+By leveraging deterministic chaos seeded from cryptographic keys and public nonces, the method enables encryption/decryption and authenticated inference based on keys without requiring retraining or alterations to original model architectures.
+The transformation operates directly on real-valued tensors in continuous space, making it compatible with modern neural pipelines and efficient to implement on hardware accelerators.
+The proposed system serves as a functional privacy mechanism: key- and nonce-conditioned, reversible, and structurally opaque to unauthorized users.
+It offers a lightweight, modular alternative to traditional cryptographic methods for use cases involving secure inference, controlled data contribution, and user-authenticated API access.
+As machine learning systems increasingly interact with sensitive and user-owned data, new methods that enable secure, interpretable, and user-aligned processing will become essential.
+This framework contributes one such approach, grounded in dynamical systems theory and designed for integration into existing model infrastructure.
+Validation experiments and empirical evaluations are ongoing.
+This preprint is shared to disseminate the concept, gather community feedback, and guide future research into learned chaotic dynamics, adversarial robustness, and secure deployment practices.
 8 Acknowledgements
-The author gratefully acknowledges the UKRI PhD studentship at the University of Edinburgh, which supported prior research training and early exploration of related ideas.
+The author gratefully acknowledges their UKRI PhD studentship at the University of Edinburgh, which supported prior research training and early exploration of related ideas.
 This work was completed independently during a period of academic leave.